OpenStudy (anonymous):

A giant pie is created in an attempt to break a world record for baking. The pie is shown below: A circle is shown with a central angle marked 42 degrees and the diameter marked 30 feet. What is the area of the slice of pie that was cut, rounded to the nearest hundredth? 78.13 ft2 82.43 ft2 86.31 ft2 91.98 ft2

OpenStudy (anonymous):

The ratio of the slice's area to the circle's area is the same as the ratio of the angle to 360° \[\frac{ 42 }{ 360 }=\frac{ x }{\pi r^2 }\]

OpenStudy (welshfella):

do you need any more help? can you work out the value of the radius ( r )?

OpenStudy (anonymous):

How do I find the radius?

OpenStudy (anonymous):

divide the diameter by 2

OpenStudy (welshfella):

radius = half of the diameter

OpenStudy (anonymous):

so id d=30 then the radius is 15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so I would do x/pi*15^2 ?

OpenStudy (anonymous):

I'd get a number for π*15² then cross multiply to find x. x is the area you're looking for

OpenStudy (anonymous):

ok so 42/360=x/706.5 then multiply 42 and 706.5 29673 so its 360x/29673 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

360x = actually

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

360x=29673

OpenStudy (anonymous):

oh ok. so its 29673 divided by the 360

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the answer would be 82.425 ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

THANK YOU SO MUCH!!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

I'm new to the site how to I give you like an awards or something?

OpenStudy (anonymous):

just hit the "best answer button" on a response from whomever you want to give a medal to

OpenStudy (anonymous):

Ok. Can you help me on one just like this one? I think I got it but I want to be sure

OpenStudy (anonymous):

A Ferris wheel car moves from point C to point D on the circle shown below: Circle A is shown with a central angle marked 38 degrees and the diameter marked 20 feet. What is the arc length the car traveled, to the nearest hundredth? 2.18 feet 4.31 feet 5.84 feet 6.63 feet

OpenStudy (anonymous):

38/360=x/pi*10^2 38/360=x/314 360x=11932 x=33.144

OpenStudy (anonymous):

Wait, I don't think I read the question right.....

OpenStudy (anonymous):

you did it right, you just used the wrong formula. Area is πr². Circumference is 2πr. Arc length is part of the circumference. 38/360=x/(2πr) 38/360=x/(2π*10) 38/360=x/62.8 360x= 2386.4 x=6.62

OpenStudy (anonymous):

Thank you!